The I/O has remained within the normal average range since 1:23PM ET. We will mark this as resolved.
An update will be provided tomorrow once the data centre investigation is complete.
Since approx. 1:23PM ET, the I/O has remained in the normal average range. We will continue to monitor closely & provide another update in 1 hour.
This appeared to be a disk network issue that was affecting the I/O performance, which is basically the disk speed incoming/outgoing. The issue started at approx. 1PM ET, resulting in sites loading slow, down & up for approx. 30 minutes.
At approx. 1:23PM ET we starting to see the I/O come back to the normal average.
Data centre staff will continue to investigate to isolate the cause & hopefully prevent the issue from reoccurring.
We're investigating what appears to be a sporadic network issue that is specific to the canada300 cluster.
Data centre staff are investigating & we will report more details here once available.